Solve the equation by a method of your choice.
x^2-2x-3=0

Found 2 solutions by rwm, algebrahouse.com:
Answer by rwm(914)

factor
what are the factors of -3 which add up to -2
-3 and 1
(x-3)*(x+1)=0
x=3 x=-1

Answer by algebrahouse.com(1659)

"Solve the equation by a method of your choice.
x^2-2x-3=0"

x² - 2x - 3 = 0
(x - 3)(x + 1) = 0 {factored into two binomials}
x - 3 = 0 or x + 1 = 0 {set each factor equal to 0}
x = 3 or x = -1 {solved each equation for x}
www.algebrahouse.com